1.Isooctane burns in air to give water and carbon dioxide (drHO=-109992 KJ/mol- rxn). What is the enthalpy change if you burn 1.00L of Isooctane?

2. The freezing point of mercury is -38.8 degrees Celsius. What quantity of energy in joules is released to the surroundings if 1.00 mL of mercury is cooled from 23.0 to -38.8 degrees celsius?
1
Expert's answer
2020-11-17T10:36:31-0500

1.) Density of isooctane= 0.69g/mL

Mass = 0.69g/mL ×1000mL = 690g

Equation : 2C8H18 + 25O2 = 16CO2 + 18H2O

2[(12×8)+ 18]g C2H18 = -109992KJ/mol

228 g C2H18= -109992KJ/mol

1 g C2H18 = -109992KJ/mol / 228g

690g C2H18 = -109992/228 × 690

= -332871 KJ/mol


2) specific heat capacity of mercury = 0.14J/g.K

Density of mercury = 5.43g/mL

Mass = 5.43g/mL ×1mL = 5.43g

H=mcΔt = 5.43 × 0.14 × (-38.8-23) = 5.43 ×0.14× -61.8

= -46.98 J